# Quasi-two-body decays $B\to Pf_{0}(500)\to P\pi^{+}\pi^{-}$ in the perturbative QCD approach

### v1

@@ I INTRODUCTION
Three-body $B$-meson decays are considerably more challenging than that of two-body-decays, mainly due to the entangled resonant and nonresonant contributions, the complex interplay between the weak and strong dynamics Charles et al. (2017), and other possible final-state-interactions (FSI) Bediaga and Magalhães (2015); Bediaga et al. (2018); Yu et al. (2024) in the three-body $B$ meson decays. Traditional approaches for the two-body-decays are no longer satisfactory in the three-body decay processes Alvarenga Nogueira et al. (2016). Practically, the hadronic three-body $B$ meson decay processes, in most cases, are considered to be dominated by the low-energy $S$-, $P$- and $D$-wave resonant states, which could be treated in the quasi-two-body framework. By neglecting the FSI between the meson pair originated from the resonant states and the bachelor particle, the factorization procedure can be applied Alvarenga Nogueira et al. (2016); Boito et al. (2017). Substantial theoretical efforts for different quasi-two-body $B$ meson decays has been made within different theoretical approaches, c.f. Refs. Gronau and Rosner (2003); Engelhard et al. (2005); Gronau and Rosner (2005); Imbeault and London (2011); Gronau (2013); Bhattacharya et al. (2013, 2014); Xu et al. (2014a, b); He et al. (2015); Furman et al. (2005); El-Bennich et al. (2006, 2009); Leitner et al. (2010); Dedonder et al. (2011); Cheng et al. (2005, 2007); Cheng and Chua (2013, 2014); Li (2014); Cheng et al. (2016); Kränkl et al. (2015); Klein et al. (2017). As well, the contributions from various intermediate resonant states for the three-body $B$-meson decays in the context of perturbative QCD (PQCD) approach Keum et al. (2001a, b); Lü et al. (2001) have been investigated in Ref. Wang and Li (2016); Ma et al. (2017a, b); Rui and Wang (2018); Li et al. (2018); Wang and Chai (2019); Ma et al. (2019); Cui et al. (2019); Wang et al. (2020); Fan and Wang (2020); Chai et al. (2022).
@@
Compared with vector and tensor mesons, the identification of the scalar mesons is long-standing puzzle Workman et al. (2022). Scalar resonances are hard to be resolved, since some of them have large decay widths which make us difficult to distinguish between resonance and background. In the theoretical point of view, their masses do not fit the expectation in the naive quark model. For the lightest scalar meson $f_{0}(500)$ (also refereed as $\sigma$) meson, which is, in the present statu, not a ordinary meson in the sense that it cannot be interpreted as predominantly made of quark and antiquark Pelaez (2016). After more than 60 years of study of $f_{0}(500)$, various of interpretations have been proposed. More explicitly, a light scalar-isoscalar field was first postulated Johnson and Teller (1955) for explanation of the inter-nucleon attraction. Then, Linear Sigma Model was proposed Gell-Mann and Levy (1960) to describe the chiral symmetry in pion-pion interaction, this explains why $f_{0}(500)$ is usually called as the $\sigma$ meson. The linear sigma model plays a relevant role, in history, for the understanding of spontaneous chiral symmetry breaking, where all fields become Goldstone bosons, i.e. pions, except $\sigma$. When the $f_{0}(500)$ and $f_{0}(980)$ are considered as $q\bar{q}$ state, the mixing of light and strange quark may appear Stone and Zhang (2013), it can be characterized by a $2\times 2$ rotation matrix with a single parameter, i.e. the mixing angle $\phi$. If the $q\bar{q}$ state does exist, the mixing angle can be constrained by the scalar decay channel which is exactly what we study. Comparison between our theoretical prediction and the experimental data may allow us to probe the inner structure of light scalar mesons. Tetraquarks is also a popular interpretation Jaffe (1977) which is bounded to form a color neutral resonance by two valence quark and two antiquark Pelaez (2016). There are also some different explanation of quark level dynamics to describe the formation of scalar mesons, such as ‘bag model’ with additional one gluon exchange Jaffe (1977), the diquark-antidiquark configurations Maiani et al. (2004), the large-$N$ quantum chromodynamics Weinberg (2013), moreover, including instanton effects ’t Hooft et al. (2008). In this theories, tetraquark states are unmixed, as assumed in Ref. Stone and Zhang (2013), with a constrain of mixing angle less than 5 degree Fleischer et al. (2011). Consequently, how to distinguish $q\bar{q}$ and $q\bar{q}q\bar{q}$ states becomes a significant issue. In Ref. Wang and Lu (2010), authors proposed a method to distinguish two kinds of scalar mesons based on sum rule technique, it may becomes a important criterion of phenomenological study and may provide more detail of inner structure of scalar mesons.
@@
Figure 1: Typical diagrams for the quasi-two-body decays $B_{(s)}\to P\sigma\to P\pi\pi$. The diagram (a) for the $B\to\sigma$ transition, and diagram (c) for the $B\to P$ transition, as well as the diagrams (b) and (d) for for annihilation contributions. The symbol $\otimes$ stands for the weak vertex and $\times$ denotes possible attachments of hard gluons.
@@
Within the frame work of PQCD approach, quasi-two body decays involving resonant state $f_{0}(500)$ have been studied Wang et al. (2015); Ma et al. (2017a); Liu et al. (2019); Xing and Xing (2019), we would like to extend to our previous studies to the quasi-two-body $B$ meson decays $B\to Pf_{0}(500)\to P\pi\pi$, where the bachelor particle $P$ denotes the light pseudoscalar $\pi$, $K$, $\eta$, or $\eta^{\prime}$. Typical diagrams for the $B\to Pf_{0}(500)\to P\pi\pi$ decays’ processes are shown in Fig. 1. Inspired by the generalized parton distribution (GPD) in hard exclusive two pion production Diehl et al. (1998); Müller et al. (1994); Polyakov (1999); Hägler et al. (2002), the two-meson distribution amplitude was introduced in three-body hadronic $B$ decays in the frame work of PQCD approach Chen and Li (2003, 2004) as the universal nonperturbative input. The decay amplitude for the quasi-two-body decays $B\to Pf_{0}(500)\to P\pi\pi$ can be expressed as the convolution of the nonperturbative wave function and hard kernel Chen and Li (2003, 2004); Wang and Li (2016)
@@
The rest of this paper are organised as follows. We give a brief introduction of the theoretical framework in Sec. II. Numerical results and some discussions are shown in Sec. III, and a brief conclusion will be summarized in Sec. IV. The relevant factorization formulas for the decay amplitudes are collected in the Appendix.
@@ II FRAMEWORK
where $\tau_{B}$ is the mean lifetime of $B$ meson, $|\vec{p}_{1}|$ and $|\vec{p}|$ are the three momenta of $f_{0}(500)$ resonance and $D$ meson respectively in the center-of-mass frame of $\pi$-$\pi$, and can be written as
$|\overrightarrow{p_{1}}|=\frac{\sqrt{\lambda(\omega^{2},m_{\pi}^{2},m^{2}_{\pi})}}{2\omega},\quad|\overrightarrow{p}|=\frac{\sqrt{\lambda(m_{B}^{2},m_{D}^{2},\omega^{2})}}{2\omega},$ (32)
with $m_{B}$, $m_{D}$ and $m_{\pi}$ are the masses of the $B$, $D$ and $\pi$ mesons respectively, and the Källén function $\lambda(a,b,c)=a^{2}+b^{2}+c^{2}-2(ab+ac+bc)$.
@@ III RESULTS
For the numerical calculation, we adopt QCD scale at $\mu=0.25$ GeV in modified minimal subtraction scheme. The decay constant of $B^{0,\pm}$, $B_{s}$ and light pseudoscalar mesons come from FLAG working group’s result Aoki et al. (2022). The masses and mean life times of $B^{0,\pm}$ and $B_{s}$ mesons, the masses of light pseudoscalar mesons and Wolfenstein parameters are all come from recent updated Review of Particle Physics Workman et al. (2022). We summarize explicitly the numerical values of the necessary standard model inputs and the hadronic parameters in Table 1.
By using the formula of differential branching fraction in Eq. (31) and explicit decay amplitude in Appendix A, we obtain the PQCD predictions of CP averaged branching fractions and direct CP violations in Table 2 for quasi-two-body $B\to Pf(500)\to P\pi\pi$ processes. The first uncertainty comes from shape parameter $\omega_{B}$ in the $B$ meson distribution amplitude, we vary it value at 10% magnitude, that is $\omega_{B^{0,\pm}}=0.40\pm 0.04$ GeV and $\omega_{B_{s}}=0.5\pm 0.05$ GeV for $B^{0,\pm}$ and $B_{s}$ mesons respectively. The second uncertainty comes from the Gegenbaur moment $a_{2}^{I=0}=0.20\pm 0.20$ in the $S$-wave $\pi$-$\pi$ distribution amplitude. We ignore the uncertainties of parameters in distribution amplitude of light pseudoscalar mesons and Wolfenstein parameters since this uncertainties are considered to be small. It is interesting to see that the $\omega_{B}$ produce the largest theoretical uncertainty in PQCD predictions of quasi-two-body $B$ meson decays Wang and Li (2016); Ma et al. (2017a, b); Rui and Wang (2018); Li et al. (2018); Wang and Chai (2019); Ma et al. (2019); Cui et al. (2019); Wang et al. (2020); Fan and Wang (2020); Chai et al. (2022), but in this study and previous quasi-two-body $B$ meson decays involving $S$-wave $\pi$-$\pi$ contributions Wang et al. (2015); Ma et al. (2017a); Liu et al. (2019); Xing and Xing (2019), the Gegenbaur moment $a_{2}^{I=0}$ plays more important role.
@@
    References       ℬ$\mathcal{B}(10^{-6})$       $\mathcal{A}_{\rm CP}(\%)$
@@
    $f_{0}(500)$ was often parameterized by Bugg model Bugg (2007) in partial wave analysis and it was applied in LHCb measurements Aaij et al. (2015a, b). However, $S$-wave $\pi$-$\pi$ are considered to be broad, overlapping resonances, $K$-matrix model was applied for parameterization of $S$-wave components as a alternative scheme Aaij et al. (2020a, b). The rigorous theoretical calculation for nonresonance contribution in the context of PQCD framework is still absent Wang et al. (2020), comparison between experiment measurements and theoretical predictions is still challenging. More attempts can be make in future study to parameterize the nonresonance contribution for sake of giving a more reliable result.
@@ IV CONCLUSION
We have studied the quasi-two-body decays $B\to Pf_{0}(500)\to P\pi^{+}\pi^{-}$ in the framework of PQCD factorization approach by using the scalar form factor $F_{\pi\pi}(\omega^{2})$ as the nonperturbative input. Here the bachelor particle $P$ denotes $\pi$, $K$, $\eta$ and $\eta^{\prime}$, respectively. The CP averaged branching fractions and the CP asymmetries for the considered quasi-two-body decay modes have been calculated by using the quasi-two-body approximation. It is found that the branching fractions are within the range of $10^{-9}-10^{-6}$. Our predicted branching fraction for $B^{+}\to\pi^{+}f_{0}(500)\to\pi^{+}\pi^{+}\pi^{-}$ agrees with the upper limit issued by the BABAR collaboration but is smaller than the LHCb measurement by a factor of ∼$\sim 2$. In Figs. (2, 3), we have shown the differential branching fraction over the $\pi$-$\pi$ invariant mass and the direct CP violation for the channel $B^{+}\to\pi^{+}f_{0}(500)\to\pi^{+}\pi^{+}\pi^{-}$. We hope all the predictions could be tested by further experiments, it may deeper our understanding for the $B$ meson decay mechanism in the context of PQCD factorization approach, and help us to probe the inner structure of $f_{0}(500)$.

### v2

@@ I INTRODUCTION
Three-body $B$-meson decays are considerably more challenging than that of two-body-decays, mainly due to the entangled resonant and nonresonant contributions, the complex interplay between the weak and strong dynamics Charles et al. (2017), and other possible final-state-interactions (FSI) Bediaga and Magalhães (2015); Bediaga et al. (2018); Yu et al. (2024) in the three-body $B$ meson decays. Traditional approaches for the two-body-decays are no longer satisfactory in the three-body decay processes Alvarenga Nogueira et al. (2016). Practically, the hadronic three-body $B$ meson decay processes, in most cases, are considered to be dominated by the low-energy $S$-, $P$- and $D$-wave resonant states, which could be treated in the quasi-two-body framework. By neglecting the FSI between the meson pair originated from the resonant states and the bachelor particle, the factorization procedure can be applied Alvarenga Nogueira et al. (2016); Boito et al. (2017). Substantial theoretical efforts for different quasi-two-body $B$ meson decays has been made within different theoretical approaches, cf. Refs. Gronau and Rosner (2003); Engelhard et al. (2005); Gronau and Rosner (2005); Imbeault and London (2011); Gronau (2013); Bhattacharya et al. (2013, 2014); Xu et al. (2014a, b); He et al. (2015); Furman et al. (2005); El-Bennich et al. (2006, 2009); Leitner et al. (2010); Dedonder et al. (2011); Cheng et al. (2005, 2007); Cheng and Chua (2013, 2014); Li (2014); Cheng et al. (2016); Kränkl et al. (2015); Klein et al. (2017). As well, the contributions from various intermediate resonant states for the three-body $B$-meson decays in the context of perturbative QCD (PQCD) approach Keum et al. (2001a, b); Lü et al. (2001) have been investigated in Refs. Wang and Li (2016); Ma et al. (2017a, b); Rui and Wang (2018); Li et al. (2018); Wang and Chai (2019); Ma et al. (2019); Cui et al. (2019); Wang et al. (2020); Fan and Wang (2020); Chai et al. (2022).
@@
Compared with vector and tensor mesons, the identification of the scalar mesons is long-standing puzzle Workman et al. (2022). Scalar resonances are hard to be resolved, since some of them have large decay widths which make us difficult to distinguish between resonance and background. In the theoretical point of view, their masses do not fit the expectation in the naive quark model. For the lightest scalar meson $f_{0}(500)$ (also refereed as $\sigma$) meson, which is, in the present state, not a ordinary meson in the sense that it cannot be interpreted as predominantly made of quark and antiquark Pelaez (2016). After more than 60 years of study of $f_{0}(500)$, various of interpretations have been proposed. More explicitly, a light scalar-isoscalar field was first postulated Johnson and Teller (1955) for explanation of the inter-nucleon attraction. Then, Linear Sigma Model was proposed Gell-Mann and Levy (1960) to describe the chiral symmetry in pion-pion interaction, this explains why $f_{0}(500)$ is usually called as the $\sigma$ meson. The linear sigma model plays a relevant role, in history, for the understanding of spontaneous chiral symmetry breaking, where all fields become Goldstone bosons, i.e., pions, except $\sigma$. When the $f_{0}(500)$ and $f_{0}(980)$ are considered as $q\bar{q}$ state, the mixing of light and strange quark may appear Stone and Zhang (2013), it can be characterized by a $2\times 2$ rotation matrix with a single parameter, i.e., the mixing angle $\phi$. If the $q\bar{q}$ state does exist, the mixing angle can be constrained by the scalar decay channel which is exactly what we study. Comparison between our theoretical prediction and the experimental data may allow us to probe the inner structure of light scalar mesons. Tetraquarks is also a popular interpretation Jaffe (1977) which is bounded to form a color neutral resonance by two valence quark and two antiquark Pelaez (2016). There are also some different explanation of quark level dynamics to describe the formation of scalar mesons, such as the bag model with additional one gluon exchange Jaffe (1977), the diquark-antidiquark configurations Maiani et al. (2004), the large-$N$ quantum chromodynamics Weinberg (2013), moreover, including instanton effects ’t Hooft et al. (2008). In this theories, tetraquark states are unmixed, as assumed in Ref. Stone and Zhang (2013), with a constrain of mixing angle less than 5 degree Fleischer et al. (2011). Consequently, how to distinguish $q\bar{q}$ and $q\bar{q}q\bar{q}$ states becomes a significant issue. In Ref. Wang and Lu (2010), authors proposed a method to distinguish two kinds of scalar mesons based on sum rule technique, it may becomes a important criterion of phenomenological study and may provide more detail of the inner structure of scalar mesons.
@@
Figure 1: Typical diagrams for the quasi-two-body decays $B_{(s)}\to P\sigma\to P\pi\pi$. The diagram (a) for the $B\to\sigma$ transition, and diagram (c) for the $B\to P$ transition, as well as the diagrams (b) and (d) for annihilation contributions. The symbol $\otimes$ stands for the weak vertex and $\times$ denotes possible attachments of hard gluons.
@@
Within the framework of PQCD approach, quasi-two body decays involving resonant state $f_{0}(500)$ have been studied Wang et al. (2015); Ma et al. (2017a); Liu et al. (2019); Xing and Xing (2019), we would like to extend to our previous studies to the quasi-two-body $B$ meson decays $B\to Pf_{0}(500)\to P\pi\pi$, where the bachelor particle $P$ denotes the light pseudoscalar $\pi$, $K$, $\eta$, or $\eta^{\prime}$. Typical diagrams for the $B\to Pf_{0}(500)\to P\pi\pi$ decay processes are shown in Fig. 1. Inspired by the generalized parton distribution (GPD) in hard exclusive two pion production Diehl et al. (1998); Müller et al. (1994); Polyakov (1999); Hägler et al. (2002), the two-meson distribution amplitude was introduced in three-body hadronic $B$ decays in the frame work of PQCD approach Chen and Li (2003, 2004) as the universal nonperturbative input. The decay amplitude for the quasi-two-body decays $B\to Pf_{0}(500)\to P\pi\pi$ can be expressed as the convolution of the nonperturbative wave function and hard kernel Chen and Li (2003, 2004); Wang and Li (2016)
@@
The rest of this paper are organized as follows. We give a brief introduction of the theoretical framework in Sec. II. Numerical results and some discussions are shown in Sec. III, and a brief conclusion will be summarized in Sec. IV. The relevant factorization formulas for the decay amplitudes are collected in the Appendix.
@@ II FRAMEWORK
where $B_{0}$ is proportional to quark condensate, in $\pi\pi$ scalar system, it can be parameterized as $B_{0}\simeq m_{\pi}^{2}/(m_{u}+m_{d})$. Following the definition from Ref. Wang (2019), the constant $C$ takes the form of $C={g_{\sigma\pi\pi}\bar{f}_{\sigma}}/(\sqrt{2}B_{0}m_{0})$. Meanwhile, $\tau_{B}$ is the mean lifetime of $B$ meson, $|\vec{p}_{1}|$ and $|\vec{p}|$ are the three momenta of $f_{0}(500)$ resonance and light pseudoscalar meson respectively in the center-of-mass frame of $\pi$-$\pi$, and can be written as
$|\overrightarrow{p_{1}}|=\frac{\sqrt{\lambda(\omega^{2},m_{\pi}^{2},m^{2}_{\pi})}}{2\omega},\quad|\overrightarrow{p}|=\frac{\sqrt{\lambda(m_{B}^{2},m_{P}^{2},\omega^{2})}}{2\omega},$ (32)
with $m_{B}$, $m_{P}$ and $m_{\pi}$ are the masses of the $B$, light pseudoscalar and $\pi$ mesons respectively, and the Källén function $\lambda(a,b,c)=a^{2}+b^{2}+c^{2}-2(ab+ac+bc)$.
@@ III RESULTS
For the numerical calculation, we adopt QCD scale at $\mu=0.25$ GeV in modified minimal subtraction scheme. The decay constant of $B^{0,\pm}$, $B_{s}$ and light pseudoscalar mesons come from FLAG working group’s result Aoki et al. (2022). The masses and mean life times of $B^{0,\pm}$ and $B_{s}$ mesons, the masses of light pseudoscalar mesons and Wolfenstein parameters are all come from recent updated Review of Particle Physics Workman et al. (2022). We summarize explicitly the numerical values of the necessary standard model inputs and the hadronic parameters in Table 1. Here the subscript $N_{f}$ in decay constant $f_{B_{(s)}}$ represents the number of dynamical quark flavor in lattice simulation. $N_{f}=2+1+1$ is for $m_{u}=m_{d}<m_{s}<m_{c}$ with four flavors dynamical quarks. The result for $N_{f}=2+1+1$ is considered to be the most realistic one in comparing with $N_{f}=2$ and $N_{f}=2+1$, which can be found by the Flavor Lattice Averaging Group in detail Aoki et al. (2022).
By using the formula of differential branching fraction in Eq. (31) and explicit decay amplitude in Appendix A, we obtain the PQCD predictions of $CP$ averaged branching fractions and direct $CP$ violations in Table 2 for quasi-two-body $B\to Pf(500)\to P\pi\pi$ processes. The first uncertainty comes from shape parameter $\omega_{B}$ in the $B$ meson distribution amplitude, we vary it value at 10% magnitude, that is $\omega_{B^{0,\pm}}=0.40\pm 0.04$ GeV and $\omega_{B_{s}}=0.5\pm 0.05$ GeV for $B^{0,\pm}$ and $B_{s}$ mesons respectively. The second uncertainty comes from the Gegenbauer moment $a_{2}^{I=0}=0.20\pm 0.20$ in the $S$-wave $\pi$-$\pi$ distribution amplitude. We ignore the uncertainties of parameters in distribution amplitude of light pseudoscalar mesons and Wolfenstein parameters since this uncertainties are considered to be small. It is interesting to see that the $\omega_{B}$ produce the largest theoretical uncertainty in PQCD predictions of quasi-two-body $B$ meson decays Wang and Li (2016); Ma et al. (2017a, b); Rui and Wang (2018); Li et al. (2018); Wang and Chai (2019); Ma et al. (2019); Cui et al. (2019); Wang et al. (2020); Fan and Wang (2020); Chai et al. (2022), but in this study and previous quasi-two-body $B$ meson decays involving $S$-wave $\pi$-$\pi$ contributions Wang et al. (2015); Ma et al. (2017a); Liu et al. (2019); Xing and Xing (2019), the Gegenbauer moment $a_{2}^{I=0}$ plays more important role.
@@
    References       ℬ$\mathcal{B}(10^{-6})$       $\mathcal{A}_{CP}(\%)$
@@
    $f_{0}(500)$ was often parameterized by Bugg model Bugg (2007) in partial wave analysis and it was applied in LHCb measurements Aaij et al. (2015a, b). However, $S$-wave $\pi$-$\pi$ are considered to be broad, overlapping resonances, $K$-matrix model was applied for parametrization of $S$-wave components as a alternative scheme Aaij et al. (2020a, b). The rigorous theoretical calculation for nonresonance contribution in the context of PQCD framework is still absent Wang et al. (2020), comparison between experiment measurements and theoretical predictions is still challenging. More attempts can be make in future study to parameterize the nonresonance contribution for sake of giving a more reliable result.
    It is also worth mentioning that the Gegenbauer moment $a_{2}$ plays very different roles in different decay mode. First, the annihilation diagram will generate very large uncertainty from Gegenbauer moment $a_{2}$. The flavor changing charged current mode such as $B^{+}\to K^{+}[\sigma\to]\pi^{+}\pi^{-}$ decay processes are transition diagram dominated, which can be seen in subdiagram (a) of Fig. 1. Meanwhile, the flavor changing neutral current mode such as $B^{0}\to K^{0}[\sigma\to]\pi^{+}\pi^{-}$ are annihilation diagram dominated, which can be seen in subdiagram (b) of Fig. 1. Taking decay process $B^{+}\to K^{+}[\sigma\to]\pi^{+}\pi^{-}$ and $B^{0}\to K^{0}[\sigma\to]\pi^{+}\pi^{-}$ as example, we can make a comparison for contributions from different diagrams. It is found that annihilation diagram will give about $72.3\%$ contribution for branching fraction in decay mode $B^{0}\to K^{0}[\sigma\to]\pi^{+}\pi^{-}$, therefore, the uncertainty generated by $a_{2}$ will very large. On the contrary, annihilation diagram will give just about $35.0\%$ contribution for branching fraction in decay mode $B^{+}\to K^{+}[\sigma\to]\pi^{+}\pi^{-}$. Thus, the large uncertainty from Gegenbauer moment $a_{2}$ will be amortized and become much lower than that of annihilation dominated mode. Subsequently, the flavor changing charged current mode are tree dominated, while the neutral mode are penguin dominated. The very complicated complex interference from CKM matrix, Wilson coefficients and form factors will enlarge or reduce the uncertainty caused by $a_{2}$, which encoding the very different behaviors in different mode.
@@ IV CONCLUSION
We have studied the quasi-two-body decays $B\to Pf_{0}(500)\to P\pi^{+}\pi^{-}$ in the framework of PQCD factorization approach by using the scalar form factor $F_{\pi\pi}(\omega^{2})$ as the nonperturbative input. Here the bachelor particle $P$ denotes $\pi$, $K$, $\eta$ and $\eta^{\prime}$. The $CP$ averaged branching fractions and the $CP$ asymmetries for the considered quasi-two-body decay modes have been calculated by using the quasi-two-body approximation. It is found that the branching fractions are within the range of $10^{-9}-10^{-6}$. Our predicted branching fraction for $B^{+}\to\pi^{+}f_{0}(500)\to\pi^{+}\pi^{+}\pi^{-}$ agrees with the upper limit issued by the $BABAR$ collaboration but is smaller than the LHCb measurement by a factor of ∼$\sim 2$. In Figs. 2 and 3, we have shown the differential branching fraction over the $\pi$-$\pi$ invariant mass and the direct $CP$ violation for the channel $B^{+}\to\pi^{+}f_{0}(500)\to\pi^{+}\pi^{+}\pi^{-}$. We hope all the predictions could be tested by further experiments, it may deeper our understanding for the $B$ meson decay mechanism in the context of PQCD factorization approach, and help us to probe the inner structure of $f_{0}(500)$.